OpenStudy (anonymous):

What is the solved systems of equations 3x+4y=16 -4x-3y=-19

pooja195 (pooja195):

Use elimintation

pooja195 (pooja195):

4 (3x+4y=16) -4x-3y=-19 Opposites cancel distribute \[\huge~4 (3x+4y=16)\]

OpenStudy (anonymous):

okay um here is the link for the first one http://www.geteasysolution.com/3x-4y=-16 and i think it will do the second one too

OpenStudy (anonymous):

its a calculator it even shows you the work it out

OpenStudy (anonymous):

just put add on it i know it says -

OpenStudy (mrnood):

@puppylife101 You have missed the point that this is a SYSTEM of equations your calculator simply re-arranges one equation - but leaves both variables (x and y) in the answer Since BOTH the equations are true together then you can solve BOTH variables from the system

OpenStudy (anonymous):

okay just helping maybe on the wrong topic opps

OpenStudy (anonymous):

I'm still confused on what the final x and y values would be

OpenStudy (mrnood):

to use the method called 'elimination' you need to get 2 equations that have one of the variables with the multiplier - e.g. BITH equations have 4x in them (for instance). To do that in THIS system I suggest this method: multiply the first equation by 4 (that will give you a term with 12x) multiply the second equation by 3 (that will give you an equation with -12x) then ADD those 2 equations and the 12x will disappear and you will have an equation with just y in it (and constants) You can then solve that equation for y If you have not seen the 'elimination method' then you should look at your text, or ask your class or study online. It is one of the basic methods for solving questions like this